OpenStudy (anonymous):

Find the real number root of radical 3 sqrt 125 over 343

OpenStudy (jdoe0001):

\(\Large { \sqrt[3]{\cfrac{125}{343}}\implies \cfrac{\sqrt[3]{125}}{\sqrt[3]{343}}\quad \begin{cases} 125\to 5\cdot 5\cdot 5\to 5^3\\ 343\to 7\cdot 7\cdot 7\to 7^3 \end{cases} }\)

OpenStudy (anonymous):

hey im confused

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

Welcome to OpenStudy! :D